40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL, Page 51 of 52  Not logged in ELOG logo
ID Date Author Type Category Subject
  92   Thu Mar 25 13:05:59 2010 FrankMiscRefCavall parts baked

by now all parts are baked and ready for assembly. The only part we couldn't bake is the chamber itself. But Bob said that putting everything together and using the heaters on the chamber to "bake" the whole thing at lower temperature would be ok. So the plan is to assemble everything including an old 8l ion pump and bake the whole thing. As soon as all the insulation, temp sensors and stuff is ready we move the cavity into the (hopefully) clean chamber and replace the 8l pump by the 20l pump currently used on the other chamber.

  91   Tue Mar 23 21:17:56 2010 Frank, TaraMiscRefCavfoam cut & partly assembled

we've cut the rest of the foam parts to fit on the smaller diameter (the main tube with the heaters). We already started to glue them together and have two halfs by now with a length of about 20". We have to make two thinner slices tomorrow to finish this inner part. We also have to glue the end caps together.

The current plan is to make 3 parts:

- one main section including the endcap parts divided into two parts which you put on the chamber from the sides

- one full endcap, which you put on the window from the end next to the edge of the table for easier access if we want to open the chamber

The current plan is to add one layer of aluminum foil and tie the three parts with aluminum tape together. We only need little force to put them together without any gap as we decided to use a 0.1" thick intermediate layer of foam which is typically used for wrapping stuff (its soft, flexible and Rod has plenty of it for packing things). This intermediate layer on top of the heater helps us to fill the gap between the heater, sensors and free space and the really stiff yellow foam. If we want to change things in vacuum we simple cut the aluminum tape along the junction and remove the endcap. This gives us full access to all the screws to open the chamber.

 

The new insulated legs should be finished by tomorrow morning...

  90   Tue Mar 23 00:26:43 2010 Frank, TaraMiscRefCavongoing work

over the weekend we baked the two AR-coated windows for the new chamber.  Bob doesn't need the oven the next couple of days so i restarted the baking again and will continue to bake the remaining parts the next couple of days. We also set up peters old vacuum pump. We got lots of stuff from 40m and cleaned all parts today. The pump is now running and pumping the whole system including the hose to the chamber. We also wrapped some heaters around the parts and started heating the stuff to make it a bit cleaner as no one knows for what the parts have been used before. They all looked pretty clean and wiping everything didn't show any obvious contamination. We can't bake it to high as some parts are viton sealed.

The new, insulated feet should be finished by wed or so. Tomorrow we start cutting the remaining foam to size and glue the parts together. I ordered the remaining heaters, one is already attached to the chamber and it fits good except that the sticky back is not sticky enough to hold it in place. The bending force is too high, so i added some aluminum tape which holds it in place now (the corners didn't want to stick). We have plenty of space for temp sensors and we will add several AD590 and a couple of platinum sensors. If we find out that the AD590 is not good enough we can easily switch to the other sensors. We should discuss how many we want and especially where. My guess is that we should add some more on this first prototype to get a feeling for the gradients or so. We can then reduce the amount on the second chamber if we want. The platinum sensors are cheap. Typically one is about $8-10 each, but  I bought a pack of 100 directly from the manufacturer and so its about $1 each only.

  89   Thu Mar 11 22:45:58 2010 Tara ChalermsongsakLaserRC noiseVCO f noise

I measured the VCO noise again.  2 methods that I tried

1) Measuring noise from a) suppressed signal (signal from pre-amp which is fed back to IFR2023) and b) error signal ((beat signal coming out of the mixer)) and match them together

Setup for 1)

ifr2023b,  carrier f = 79.994620 Hz, power =    7.0 dBm,     FMdevn = 500kHz

mixer     Mini Circuit ZFM-3-S+   (Lo input = 7dBm, RF input = -1.05 dBm)

pre amp, SR560, Gain inv10, low pass at 1Hz

 

2) Lowering the unity gain frequency and measuring only the error signal (beat signal coming out of the mixer)

Setup for 2)

ifr2023b,  carrier f = 79.995 316 Hz, power =    7.0 dBm,     FMdevn = 1kHz

mixer     Mini Circuit ZFM-3-S+   (Lo input = 7dBm, RF input = -1.05 dBm)

pre amp, SR560, Gain inv20, low pass at 0.03Hz

 

Converting V/rHz to f/ rHz

1)The feed back signal can be convert to freq noise by using the calibration from last week, which is 0.35 MHz/V, since two setups on IFR2023 are similar.

We can obtain the calibration by giving input voltage to the LO and see how the carrier freq changes.

 

2) For the  error signal out of the mixer, disconnect the feedback signal and measure the slope of the signal output. That will be [V/rad] calibration.

Then we can convert V/rHz -> rad / rHz , multiply by the corresponding f to get f/rHz

 Red and Green plot are from the 1st method,

Blue plot if from the 2nd method.

The results between two methods are not even close to each other, I'll check tomorrow to see if I did something wrong,

Attachment 1: VCO_noise.png
VCO_noise.png
  88   Fri Mar 5 14:40:21 2010 FrankLaserRC noiseNoise from VCO

Quote:

Quote:

VCO frequency noise is measured. V input is 4.7 Volt. The signal output from the VCO (which controls the AOM) is mixed with signal from ifr2023 at ~80MHz. The demodulated signal

is then fed to SR560. Gain setting is 5 x10^3, low pass at 1Hz. The output signal is split into two. One is sent back to ifr2023 for freq modulation, another one is fed to the spectrum analyzer. The voltage output is

then converted to frequency by calibration from last week (0.714 MHz/V).

The plot shows frequncy noise from:

1: function generator (ifr2023) in brown

2: VCO,  in pink

3&4 Frank's RC noise data from FEB09, in red and blue

5:  RCnoise from beat measurment, in green

6: estimated noise

 

VCO seems to limit our noise at f=100Hz and higher/

 you have to measure a new calibration coeff. If you change the center frequency of the marconi the maximum FM modulation range changes too and so the coeff. If i remember right your range was 500kHz compared to 1Mhz last week, right? So it should change about a factor of 2 or so, so the actual measured noise is a little bit less but nevertheless much to high

 Yes, I completely forgot about that. I calibrated it again, at 80Mhz, 7.0 dBm, freq devn 500khz. and it is 0.35655 MHz/volt. About a factor of 2 smaller.

I'll put up the corrected plot soon.

Attachment 1: irf2023_@80Mhz.png
irf2023_@80Mhz.png
  87   Fri Mar 5 00:34:09 2010 FrankLaserRC noiseNoise from VCO

Quote:

VCO frequency noise is measured. V input is 4.7 Volt. The signal output from the VCO (which controls the AOM) is mixed with signal from ifr2023 at ~80MHz. The demodulated signal

is then fed to SR560. Gain setting is 5 x10^3, low pass at 1Hz. The output signal is split into two. One is sent back to ifr2023 for freq modulation, another one is fed to the spectrum analyzer. The voltage output is

then converted to frequency by calibration from last week (0.714 MHz/V).

The plot shows frequncy noise from:

1: function generator (ifr2023) in brown

2: VCO,  in pink

3&4 Frank's RC noise data from FEB09, in red and blue

5:  RCnoise from beat measurment, in green

6: estimated noise

 

VCO seems to limit our noise at f=100Hz and higher/

 you have to measure a new calibration coeff. If you change the center frequency of the marconi the maximum FM modulation range changes too and so the coeff. If i remember right your range was 500kHz compared to 1Mhz last week, right? So it should change about a factor of 2 or so, so the actual measured noise is a little bit less but nevertheless much to high

  86   Thu Mar 4 19:44:08 2010 Tara ChalermsongsakLaserRC noiseNoise from VCO

VCO frequency noise is measured. V input is 4.7 Volt. The signal output from the VCO (which controls the AOM) is mixed with signal from ifr2023 at ~80MHz. The demodulated signal

is then fed to SR560. Gain setting is 5 x10^3, low pass at 1Hz. The output signal is split into two. One is sent back to ifr2023 for freq modulation, another one is fed to the spectrum analyzer. The voltage output is

then converted to frequency by calibration from last week (0.714 MHz/V).

The plot shows frequncy noise from:

1: function generator (ifr2023) in brown

2: VCO,  in pink

3&4 Frank's RC noise data from FEB09, in red and blue

5:  RCnoise from beat measurment, in green

6: estimated noise

 

VCO seems to limit our noise at f=100Hz and higher/

Attachment 1: Noise20100304.jpg
Noise20100304.jpg
  85   Wed Mar 3 15:59:10 2010 Tara ChalermsongsakLaserRC noiseNoise from function generator

 I took the data of frequency noise of the functiongenerator("Marconi") and spectrum analyzer's noise from Mott's elog on  Nov 13 14:37:11 2009,  AdhikariLab.  Thanks, Mott

 To measure the noise @160Hz, two Functiongenerators are set at 160MHz, then mix two signals togather to get phase noise. Multplying the phase noise by corresponding frequencies to get frequency noise. 

Assuming two ideal function generators, the freq noise is divided by sqrt(2) to get noise contributed by one function generator.

The attached graph shows frequency noise from Marconi noise, detector noise(spectrum analyzer that measured Marconi noise) , beat noise( noise from beating two beams after the cavities), and estimated noise.

VCO noise will be updated soon.

Attachment 1: Noise.pdf
Noise.pdf
Attachment 2: Noise.fig
  84   Wed Mar 3 08:46:26 2010 ranaLaserRC noiseNoise as measured by the 40m MC

This link points to a measurement of the frequency noise at the 40m.

'MC_F' means the feedback to the VCO used to keep the MC locked. Below ~100 Hz, there is also feedback to the MC length and so you cannot assume its frequency noise.

From 100-1000 Hz its all acoustic pickup on the PSL table. Above 1 kHz, I believe its all VCO phase noise.

I also suggest that from now we all decide to post the actual data along with the plots we post here. Also the .m matlab file which generates the plot as done at GEO. This makes it much easier to reproduce the result a year down the line.

  83   Mon Mar 1 22:39:39 2010 ranaLaserRC noiseNoise Budget

40m entry on RIN induced thermo-optic noise here

  82   Mon Mar 1 17:52:46 2010 Tara ChalermsongsakLaserRC noiseNew Result

I got a quarter wave plate from Greg Ogin this afternoon. The attached plots show:

black-> RC noise when there are no quarter wave plates. It's the beat of circularly pol beams.

green-> when one wave plate is intalled.(from last week)

Blue-> two wave plates are in used. I'm surprised that nothing changed much from green.

Red-> I use the Buzby box to amplifie the demodulated signal and connect it to SR560 for filtering. The SR560 complains about overloading signal when I

set gain at 1, but it's ok when gain is 2.

It seems still far from noise budget Rana gave me. The peak around 1000 Hz might to be investigated.

Attachment 1: Mar01_RCnoise.png
Mar01_RCnoise.png
  81   Sun Feb 28 23:53:21 2010 ranaLaserRC noiseNoise Budget

rcnb.png

Check out the PSL directory from the SVN to see the details of the calculation.

  80   Fri Feb 26 15:15:56 2010 Tara ChalermsongsakLaserRC noisequarter wave plate added/power from Ref and A Cavs measured

I borrowed one of the quarter wave plate and added it after the beam from the Ref Cav. One more QWP is needed for the ACav.

The power after the Ref cav is measured:

1) just after the cavity : 5.4 mW

2)First BS: Reflected beam (to the PD): 3.8 mW, Transmitted beam: 1.4 mW

3) 2nd BS: Reflected beam (for beat measurement): 0.56 mW,  Transmitted beam(to the camera): 0.62 mW

The waveplate is set at 32 degree, for max transmitted beam from a PBS( horizontal polarization in this setup)

 

Power after the A cav:

1) just after the cavity: 19.3 mW

2) First BS: Reflected beam(to PD):19 mW, Transmitted beam: 0.21 mW

3) 2nd BS: Reflected beam(for beat msmt): 0.255 mW

  79   Fri Feb 26 11:40:47 2010 FrankComputingDAQacav VME crate stopped working

last night one of the DAQ cards failed and the acav crate stopped working, so also the temp stabilization of the analyzer cavity stopped woking. I restarted everything this morning and the setpoint should be reached again by lunch time or so

  78   Thu Feb 25 19:03:59 2010 Tara ChalermsongsakLaserRC noiseNoise floor from PD and SR560

Noise floor from PD and SR560 are measured, then converted from V/rHz to frequncy/rHz by the calibration from IFR2023 (0.7 MHz/Volt).

 SR560 has the same setting (gain invert 100, low pass filter at 1 Hz) as it did yesterday (when I measured the power spectrum

of the feed back signal.) The power spectrum of the feedback signal (from yesterday) is plotted in gray,  noise floor from PD( mixed with 160MHz signal from IFR2023) is plotted in pink, and noise from SR560 is in blue. I'll get the RC noise data on 2010Feb09, so I can plot them on the same graph.

Attachment 1: noise_floor.png
noise_floor.png
  77   Wed Feb 24 20:07:00 2010 Tara ChalermsongsakLaserRC noiseRC frequency noise

 I changed the gain setting on SR560 to find how it will effect the noise floor of the RC noise.

It seems that the gain and the cutoff frequency do not alter anything at lower frequency (below 100Hz), but they

change the position of the peak around 1kHz.

I'll find out noise from SR560 and the PD to see if their noises are dominating in this f.

Attachment 1: rf_noise_feb24_2010.png
rf_noise_feb24_2010.png
  76   Wed Feb 24 19:40:25 2010 Tara ChalermsongsakLaserRC noisecalibrating ifr2023

The calibration for ifr203 input for external frequency modulation is 0.714 MHz/Volt. For 160 MHz carrier, @7dBm, over +/- 1V range.

Attachment 1: calibration_feb_24.png
calibration_feb_24.png
  75   Wed Feb 24 11:58:20 2010 ranaLaserRefCavsuccessfully use ifr2023 to track the frequncy drift of the beat

Volts? What are Volts???

This plot should be converted into radians/rHz or Hz/rHz in order to be used.

Plots without physical units should almost never be used. Always Calibrate.

  74   Tue Feb 23 20:13:00 2010 Tara ChalermsongsakLaserRefCavsuccessfully use ifr2023 to track the frequncy drift of the beat

  The drift in the demodulated signal (beat frequency x local oscillator) can be tracked by using ifr2023 and sr560. We successfully set the control loop, but

the detail about how ifr2023 works will be reviewed for clarity. The SR560 is set at gain -100, low pass at f = 10Hz. The power spectrum of the drift can be seen from the attachment.

Attachment 1: 02_23_2010.pdf
02_23_2010.pdf
  73   Mon Feb 22 18:58:23 2010 Tara ChalermsongsakLaserRefCavbeat frequency between transmitted beams

 Today  we finally see the beat between transmited beams from Ref Cavity and A Cavity.  Now we are trying to use a local oscillator [IFR2023b] to demodulate the signal.

The beat signal will be fed to a low pass amplifier and  sent to IFR2023 as external  frequency modulation.

Attachment 1: 02.TIF
02.TIF
  72   Mon Feb 22 16:21:43 2010 Frank, TaraLaserRC noiseboth cavities locked, optimizing beat signal

tuned the temperatures over night to be able to lock both cavities today. Cavities are now locked and Tara is optimizing the beat signal.

  71   Mon Feb 22 11:25:51 2010 FrankLaserLaserlaser fault again

the laser stopped working while beeing in the lab but not touching the laser or table or anything. I was looking for an EOM mount and suddenly the chiller beeped. The power supply stopped working with a "PS error" message, but the more interesting fact is that the setpoint of the chiller was changed to 15degC (my guess is from the PS), but it should be at 26.5degC. So maybe the PS is also causing the freezing of the chiller, simply sending the wrong setpoint values when having a fault...

  70   Fri Feb 19 18:23:24 2010 Tara ChalermsongsakLaserRefCavalignment for measuring beat between 2 cavities

I finished aligning the beams from Ref cavity and Acavity, broke a connector on the PD we were going to used to measure the beat .

Frank borrowed another PD to use for now and ordered the replacement for the broken connector which should arrive next week.

Now I'm waiting for the temperature to settle, so both cavities can be locked and see the beat.

  69   Fri Feb 19 00:31:40 2010 FrankLaserLaser100mW lightwave NPRO slope

Peter gave me the hint that the default values are stored in an eeprom in the laser head. So connecting the head to a driver not used before shows the default values for the head. For this head it's 0.86A. So i measured the slope of the NPRO up to that value, reaching the 100mW at the default value stored in the head without tuning the diode temperature. So the head seems to be refurbished and not dead...

NPRO_slope_2010-02-19.png

  68   Thu Feb 18 15:07:39 2010 FrankMiscLaserboth IR viewers missing

so please bring it back - we are waiting for it and can't go on without those... For the future: if you take it, bring it back and don't wait until someone is missing it...

  67   Wed Feb 17 15:40:04 2010 FrankLaserLaserold 100mW NPRO

does anyone know the typical operating current for the 100mW lightwave laser model ? (M126N-1064-100) It's typically ~1.1A for the 200mW model. I've set up everything and it starts to lase around 0.44A, so at least its not dead but i don't know how high up i can go. My guess is that it is something around 0.8A but i have no datasheet which tells me...

  66   Tue Feb 16 19:33:53 2010 FrankLaserLaserseveral shutdowns the last days

we had several shutdowns of the laser within the last days. A couple of times the well known "HT error", today we had an "PS error" for the first time. When does this happen? The other error is related to a malfunction of the chiller as we found out by luck. The chiller temp readout jumps from 26 down to 15 or so within a fraction of a second (so it's not real). This causes the PS to start heating even if the temp is high enough. This screws up the stabiliy of the laser and sometimes causes a chiller error as well. But the "PS error" ? Any idea?

  65   Fri Feb 12 12:46:38 2010 Tara ChalermsongsakLaserLasera picture of beam path on the optic bench

I took a picture of the setup in PSL lab, and drew a line for laser path. I omit the mode cleaning part since it's not in use now.

 

Attachment 1: psl_full_beam.jpg
psl_full_beam.jpg
  64   Wed Feb 10 18:04:01 2010 Tara ChalermsongsakLaserRefCavQPD calibration

Today I calibrated the QPD on the ref cavity. The armlength from AOM to the quad is 2.1 m.

The calibration for x and y channel on the QPD in radian unit

QPDX:  radian = [dV]* 10^-3 / (3.36 * Vo)

QPDY: radian = [dV] * 10^-3 / (3.689 * Vo)

where [dV] is reaout voltage in Volt, Vo is the sum voltage from 4 quadrants in Volt.

 

If the prefer units are in micro radian and milli volt

QPDX:  urad = [d_mV] / (3.36 * Vo)

QPDY: urad = [d_mV]/ (3.689 * Vo)

where [d_mV] is reaout voltage in milliVolt, Vo is the sum voltage from 4 quadrants in Volt.

  63   Wed Feb 10 04:01:50 2010 ranaLaserRC noiseeven newer results

Looking better. I'm curious about what the existing loop shapes are. The old FSS hardware is designed to drive an NPRO + 1 EOM. Is that the existing layout?

However, I think its not designed to drive an EOM. The PDH box should be able to drive the VCO if we replace the output OP27 with a TLE2027. The main point is that the refcav only provides a pole @ 40 kHz and we need the electronics to be 1/f below there. The old FSS board used to do this by the combination of the series resistor and the NPRO PZT capacitance. The VCO is not a capacitive load. I guess that Tara is working on a Simulink model of this whole setup?

I kind of doubt that we will have success without using resonant RFPDs. In most of the PSLs we use a pretty large modulation depth and its necessary to really tune the 2f trap to get rid of the J1^2 term.

------

Different idea: why not just use the VCO/AOM to lock to the ACAV? Then if we pick off the beam for the RCAV after the AOM, the feedback to the NPRO can be used as the differential cavity signal. In this way, we are not sensitive to the VCO calibration issues since its squashed by the ACAV loop.

Idea #2: Just use the transmitted light from the cavities and beat them. Its only phase detection, but in principle, this is easily good enough to detect what we want. Also RF sidebands are gone and all we need is an 1811 or such to detect the beat signal.

  62   Tue Feb 9 18:33:59 2010 FrankLaserRC noiseeven newer results

measured a couple of times today with everything re-aligned and different gain settings for the FSS stuff. Measured also to lower frequencies. The problem here is that the frequency band of interest can not be measured with epics (to slow) and the measurement using the SR785 takes so long that the operating point of the VCO and therefore the coefficient changes during the measurement and different spectra don't fit together quite well. so we have to measure a couple of times to get some measurements fitting together when the changes are so small that we don't see it with our eyes. this gives us a new upper limit of:

refcav-noise_2010-02-09.png

maximum gain settings:
CG : 24dB
FG : 13dB

  61   Mon Feb 8 22:03:32 2010 FrankLaserRC noisenew results

everything is back, pointing is much much less, power modulation too, ugf of the other loop is much higher - but nevertheless the performance is not much better, only one order of magnitude. it turns out that the noise is limited by the existing FSS stuff. tuning the gains (common, fast) can produce almost any shape and level (see graph), except much lower levels are not possible, at least not for more than a couple of seconds before everything starts to oscillate. so i will to debug the old FSS stuff first to see whats going on there. will try to investigate the noise of the FSS loop and maybe replace the VCO by a function generator and probably the RF photodiode to see if that changes something. Those can be exchanged easily without changing to much...

refcav-noise_2010-02-08.png

  60   Mon Feb 8 18:27:42 2010 Tara ChalermsongsakLaserRefCavChaging Isomet AOM to Crystal Tech AOM

After verifying that the Isomet AOM causes the power modulation and realigning it did not seem to help improving the pointing effect, we switched to a Crystal Tech AOM. The 1st order diffracted beam has 55% of the total power, the remaining power before entering the cavity is ~15mW. The pointing effects observed from QPD improved a little bit to 500 mV on X and Y directions.

  59   Mon Feb 8 18:12:46 2010 FrankNotesAOMimpedance measurements

measured the impedance from 70MHz to 90MHz for three different AOMs we had in the lab - two Isomet 1205C-843 and one Crystal Technologies 3080-194.
It turns out that the tiny power modulation on top of the big power modulation and pointing effects is related to some resonance effects in the AOM.
The Isomet AOMs show both this feature, the other one not. So we decided to install the one from Crystal Technology...

AOM-impedance_2010-02-08.png

  58   Mon Feb 8 18:08:53 2010 FrankMiscAOMthermal images of AOM

took some images with the thermal imaging camera of the AOM installed in the PSL so far. The first three pictures show the AOM driven with 75MHz, 80MHz and 85MHz.

85MHz:
IR20100208_0013.jpg

80MHz:

IR20100208_0014.jpg

75MHz:
IR20100208_0012.jpg

it is interesting that the hottest zone is at the end of the crystal, not where the pzt is mounted. It looks like the crystal is not proper mounted. Here a normal image for comparison...
P1410925.JPG

so i took some pictures of a different AOM, but same model. Here are the pictures for different power levels @ 80MHz:

IR20100208_0029.jpg

IR20100208_0030.jpg

IR20100208_0024.jpg

this looks normal...

*** Rana: I've replaced Frank's AOM picture with a zoomed in one. This circuit looks a little sloppy to me - why are the coils so loose?

Attachment 4: P1410925.JPG
P1410925.JPG
Attachment 7: IR20100208_0024.jpg
IR20100208_0024.jpg
  57   Sun Feb 7 20:56:03 2010 FrankLaserRefCavAOM power modulation & pointing

seems like a bad AOM and an impedance matching problem together. exchanging the semi-rigid cable by other cables produces a lot of different results in power modulation / pointing vs frequency. right now i found a combination where for frequencies 80MHz and higher it is almost flat, but below it drops a lot. the pointing is related to the absorbed power in the AOM. - you can really feel the heat of the crystal, no joke. will measure the temp on tomorrow for different frequencies where the pointing is worst case to see if it is a macroscopic change in temp. Aidan has this nice thermal image camera to do this. i think we should try the crystal technology AOM we have. i don't think that aligning helps a lot here. tried this the whole afternoon :-(  the lowest power modulation over the entire frequency range is about 20%pp

 

  56   Sun Feb 7 18:29:06 2010 FrankLaserLasernew channels + changes

added the following channels:

C3:PSL-RCAV_DIFFPWR : diffracted power (single pass) measured behind curved mirror
C3:PSL-126MOPA_PWRMON : laser output power monitor measured after PC

 

changes:

C3:PSL-RCAV_QPDSUM changed back to QPD sum signal

 

all signals available in both framebuilders

  55   Sat Feb 6 12:19:14 2010 ranaLaserRefCavstrange things going on

Seem like a screwy AOM. I would take the double-passed beam and beat it with the initial beam (no cavities). Make sure the frequency shift is appropriate and make sure there is no amplitude change in the beat over the whole VCO range.  

  54   Fri Feb 5 21:53:13 2010 FrankLaserRefCavstrange things going on

in order to get an idea about the part/subsystem causing the RF power modulation while sweeping the VCO frequency i changed the length of the cable to the AOM. I decided to just extend the existing cable using a cheap SMA cable, not one of those semi-rigid ones. Now i thought due to more losses, low quality cable, longer, more connection etc. the shifted power should be less, but it seems to be the opposite, see screenshot. The purple curve is the transmitted power through the refcav, red the VCO input signal, the rest unimportant. On the left third the old cable, then a short break where i extended the cable, the center part with the longer cable, another short break and then with the original length. First i thought it's some aligment related thing, touching the cable, changing the alignment of the AOM etc. But it's not. The PD behind the curved mirror (not shown) shows exactly the same. Any ideas?

 

AOM-cable-length.JPG

  53   Fri Feb 5 21:10:48 2010 FrankLaserRefCavnew PD for frequency-shifted power detection

added a detector behind the curved mirror to measure the changes in deflected power (1.order) while changing the frequency of the VCO. PD is a large area Si diode to reduce pointing effects. Signal is connected to QPD-SUM instead of the QPD signal (QPD channels for x and y still connected and valid, only sum disconnected !!). We didn't trust the sum anymore as we can see slightly pointing related changes there as well.

  52   Fri Feb 5 20:53:58 2010 FrankLaserRefCavlatest values for refcav

values for  VCO pwr 4.7V

RCTRANSPD signal: 5.17V

DC-level of RF-PD:
refcav unlocked: 590mV
refcav locked : 202mV

current cavity temp: 67.7degC
slow actuator value : 0.8966V

  51   Fri Feb 5 20:37:06 2010 Tara ChalermsongsakLaserRefCavbetter pointing alignment

 After adjusting the mirror for reflecting beam back to the AOM, the QPD signal shows the better alignment. Before the voltage different readout is about 1.2 V, now it's reduced to ~500 mV.

We'll try to add a translational stage for the mirror for better alignment.

  50   Fri Feb 5 05:13:09 2010 Peter, FrankLaserRefCaveverything back online

we spent the whole night to re-align everything. By now everything is back online, both cavities were locked.for a short time. We still have the pointing problem. It's different now, seems to be smaller but is now almost the same for x and y. Will do this later today ....

The 10kHz peak in the old spectrum seems to be related to the crossover frequency of the fast actuator and the pc of the fss loop.

The PDH box is modified, all filter stages now have socket adapter boards to easily change the filter frequencies by changing the capacitor value within seconds to optimize everything. A zero for compensating the cavity pole is also installed.  A modulation input too.

  49   Thu Feb 4 22:01:27 2010 FrankLaserRefCavacav re-aligned

acav is now re-aligned. As Tara stopped aligning the AOM i will use the NPRO pzt to lock it (instead of the VCO). Then i will also see where the shit of the other tf comes from (hopefully).

  48   Thu Feb 4 20:13:53 2010 FrankLaserLaseroutput coupler to acav changed

exchanged the old mirror (T330-HR, T331-AR) by a simple Y1-1025-45P to get more power.

measured laser power : 7.17W
downstream of the new output coupler : 134.6mW

added waveplates & pbs to make the power adjustable. current power through the EOM is 8mW which gives about 4.33V on the RF-PD (Thorlabs PDA10CS, 0dB-setting, 17MHz)

  47   Wed Feb 3 23:41:11 2010 FrankLaserRefCavTF of ACAV (VCO+cavity+pd+mixer)

measured tf of the analyzer cavity including VCO (wideband input, around 80MHz center frequency), double-passed AOM, cavity, pd (Thorlabs PDA10CS) and mixer for different source levels (to see if there are nonlinearities).

TF_acav_pdh_all.pdf

  46   Wed Feb 3 18:58:50 2010 TaraLaserRefCavfrequency findings

When the VCO is modulated by a triangular function +/- 6V, a small fluctuation is observed in transmitted and reflected beam. There might be some small constant phase shift between the two.

at 78 MHz  the modulation width is about 0.3 MHz, with +/- 2.4% amplitude

 at   79.5 MHz,  the width is 0.238 MHz, +/- 1.1 % in amplitude,

  at 81 MHz, the width is 0.193 MHz, +/- 1.5% in amplitude.

The first plot shows the comparison between the transmitted beam and the reflected beam between 78 - 82 MHz. The second and the third plots are the whole data from the transmitted beam and the reflected beam respectilvely. The data is taken aroud 2010-02-03 20:00:00 UTC.

Attachment 1: comparison.JPG
comparison.JPG
Attachment 2: Tbeam.JPG
Tbeam.JPG
Attachment 3: Rbeam.JPG
Rbeam.JPG
  45   Wed Feb 3 16:19:52 2010 Tara ChalermsongsakLaserRefCavfrequency findings

Quote:

Quote:

I guess if you have a frequency counter with a GPIB interface or a simple flip-flop XOR based phase/frequency discriminator, you can feed the output to a 3113 and use at as an input to a slow EPICS PID to bring the beat frequency to within range.

Actually, we need a frequency discriminator for the Green Locking so it might be good to brainstorm about this with Aidan and Koji.

 good idea -  we now know where we have to be +- acouple of 100mK as the impact of changes in room temp is not equal for both systems so we have to slightly adjust the temperature of one cavity.

Right now we have a much larger problem with the AOM. if we lock both cavities and the AOM frequency changes slowly over time (both cavities drift a bit) we get a huge amplitude modulation for the refcav, meaning a drift in power from almost 0 to 2.5V (the full range) topped by a sine-like modulation, looks like an etalon with about 10% modulation depth, every 1.4Mhz. It's not the power of the VCO, that changes only by about .2%, but you see this modulation there as well, but almost covered by the noise of the DAQ already. It's a power modulation of the diffracted beam as you can see it in the reflected and transmitted light with the same sign, so it's not pointing or so. I think it could be an impedance matching problem causing some standing wave resonator. We try to investigate it today. I will post a graph later today...

 The power at the AOM and the reflected beam from the AOM  vs frequency are measured, but they seems to be unrelated to the problems.

Attachment 1: power_at_AOM.jpg
power_at_AOM.jpg
Attachment 2: Reflected_pwr_from_AOM.jpg
Reflected_pwr_from_AOM.jpg
  44   Wed Feb 3 13:05:26 2010 FrankLaserRefCavbeam pointing @ refvac

measured the beam pointing caused by driving the AOM frequency modulation input. data is uncalibrated so far, just a screenshot of the dataviewer. PDHOUT is the VCO input signal...

pointing.jpg

  43   Tue Feb 2 18:22:38 2010 FrankElectronicsDAQQPD channels

QPD channels for RefCav beam pointing measurements:

C3:PSL-RCAV_QPDX : X

C3:PSL-RCAV_QPDY : Y

C3:PSL-RCAV_QPDSUM : SUM

ELOG V3.1.3-